OpenStudy (tinash):

Solve 3(x+3)/4 - 2(x-3)/3 = x+2/2

OpenStudy (jiteshmeghwal9):

\[\frac{3(x+3)}{4}-\frac{2(x-3)}{3}=\frac{x+2}{2}\]Taking LCM of 3 & 4\[{9(x+3)-8(x-3)\over12}={x+2 \over 2}\]

OpenStudy (jiteshmeghwal9):

simplifying further\[{x+51 \over 12}={x+2 \over 2}\]cross multiplying\[2(x+51)=12(x+2)\]

OpenStudy (jiteshmeghwal9):

now solve for x

OpenStudy (tinash):

Would that be 2x+102-12x+24